15 moles A, 18 moles of B and 20 moles of C reacted together according to the reaction A + 2B + 4C  AB2C4. The moles of AB2C4 formed is
 

Grade:

1 Answers

Pratyugna Manna
5 years ago
As one can see both A and B are the excess reagent and C is the limiting reagent , we will apply POAC principle of C atoms because it remains conserved in the reaction. Remember that you cannot apply POAC on excess reagents.
So we  write as 
1*(moles of C)=4 * (moles of AB2C4 formed)
=>  20=4*(moles formed)
=> Moles of AB2C4 formed is 5 moles